The answer must be given In radians. Thanks!
2cos^2x + 3cosx -2 = 0

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

The answer must be given In radians. Thanks!
2cos^2x + 3cosx -2 = 0
----
Factor::
(2cos(x)-1)(cos(x)+2) = 0
------
cos(x) = 1/2 or cos(x) = -2 (extraneous)
-----
x = arccos(1/2) = pi/3 or x = (11/3)pi
----------------------------------------------
Cheers,
Stan H.
-----------------

Answer by ikleyn(52898)

.
The answer must be given In radians. Thanks!
2cos^2x + 3cosx -2 = 0
----
Factor::
(2cos(x)-1)(cos(x)+2) = 0
------
cos(x) = 1/2 or cos(x) = -2 (extraneous)
-----
x = arccos(1/2) = pi/3 or x = (5/3)pi